OpenStudy (anonymous):

convert 3/8 into percent

OpenStudy (dumbcow):

divide to convert to decimal ... then move decimal 2 places to right to make it a percent

OpenStudy (dumbcow):

another method uses proportion: \[\frac{P}{100} = \frac{3}{8}\] \[P = \frac{300}{8}\]

OpenStudy (anonymous):

How I am a 60 year old lady and do not know how to do it.

OpenStudy (dumbcow):

haha well calculator is your friend :) some even have a percent button

OpenStudy (anonymous):

Yes but I need a class in learning to use the stupid calculator. I could never understand this stuff in high school and now its just one of those credit fillers in order to get my associates degree in Medical coding and Billing. I hate fractions and I especially hate algebra. I could learn Spanish easier and standing on my head at the same time.

OpenStudy (anonymous):

I have one more problem and it involves converting a decimal into a percent. Are you up for another challenge?

OpenStudy (dumbcow):

sure